Solve the equation for x and check.
(3x-5)/18+x=(3x+5)/3+1
can you please show work

Answer by Boreal(15235)

The position of the parentheses matter.
[(3x-5)/18] +(x)=[(3x+5)/3] +(1)
Put everything over a common denominator of 18
(1/18){(3x-5+18x)}=(1/18){(6(3x+5)+18}
We can remove the (1/18), since it is common to both, and then combine terms.
21x-5=18x+30+18
3x=53
x=53/3
check
(53-5)/18 +(53/3)=(48/18)+(318/18)=-(366)/18=61/3
Right side: (53+5)/3 +(3/3)=61/3
If the parentheses are different, such as (3x-5)/(18+x), then there will be a different solution.